OpenStudy (anonymous):

ABC has vertices A(0,6), B(4,6) and C(1,3). Find the orthocenter. Will give medal

OpenStudy (anonymous):

@dan815

OpenStudy (dschneider2016):

y=-1x+z 6=-1*0+z z=6 so y=-1x+6 So we have 2 lines x=1 and y=-1x+6 They intersect at (1,5)

OpenStudy (dschneider2016):

so (1,5) is your answer